[Hallc_running] [New Logentry] Status Update

pooser at jlab.org pooser at jlab.org
Wed Feb 28 12:20:01 EST 2018

Logentry Text:
MCC completed RF recovery at 12:45 pm yesterday and beam was restored to the Hall.  We commenced with collecting production data with the LH2 target while the SHMS was at -2.0 GeV and 39 degrees and the HMS was at -1.5 GeV and 59 degrees.  The RF recovery proved to quite be beneficial since the accelerator was able to deliver stable 22 uA to Hall A and 50 uA to Hall-C (https://logbooks.jlab.org/entry/3539791) which was a ~20% increase for Hall-C.

At 5:52 pm MCC informed us that our fast raster failed (https://logbooks.jlab.org/entry/3539804).  Luckily, Bill G. was close to the lab and was able to make an escorted access and replace the failed 24 V power supply for the beam side y magnet (https://logbooks.jlab.org/entry/3539828).  This work was completed just inside 1 hr so muchos kudos to Bill G. for getting us back up and running so quickly!  The fast raster distribution was verified later be functioning nominally (https://logbooks.jlab.org/entry/3539885).

At ~8:20 pm we completed the HMS kinematic point at -1.5 GeV and 59 degrees.  The HMS momentum was then decreased to -1.35 GeV and collection of production data commenced with LH2.  For the remainder of swing and owl, runs were taken with 2 hrs on LH2, 1 hr on 10 cm Al dummy, followed by 2 hrs on LD2.  Throughout the night we had relatively smooth running (https://logbooks.jlab.org/entry/3540131).  In fact, there was 1 hour of 65 uA on LD2 target as a result of Hall A being down as a result of laser issues.  For the SHMS kinematic point of -2.0 GeV and 39 degrees we have collected 100% of the LH2 data, 63% of the LD2 data, and 56% of the 10 cm Al dummy data.  If we maintain relatively stable 50 uA beam we will likely move to the next kinematic point of -1.6 GeV while staying parked at 39 degrees sometime during swing shift tonight.  For the HMS we have collected 40% of LH2 data, 44% of LD2, and ~70% of 10 cm Al dummy.

Currently there are no scheduled accesses planned however, if an access does present itself we will contact Tony Delacruz to swap out the CW dump viewer camera since it has recently died.

Special notices in effect for shift crews:
Recovering the HMS DAQ: https://hallcweb.jlab.org/wiki/index.php/Hall-C_HMS_DAQ_Failure_Modes
Trouble shooting HMS Q1 valves JT6 and JT8: https://logbooks.jlab.org/entry/3538239, https://logbooks.jlab.org/entry/3538243, https://logbooks.jlab.org/entry/3538254
Trouble shooting HMS Q2 power supply: https://logbooks.jlab.org/entry/3537306
TO Instructions for LH2 <--> LD2 switches : https://logbooks.jlab.org/entry/3534487

We will meet today in 200C at 2:20 pm (second floor of the counting house)
Link: https://bluejeans.com/543903672
Connecting directly from a room system?
1) Dial: or bjn.vc
2) Enter Meeting ID: 543 903 672
Just want to dial in on your phone?
1) Direct-dial with my iPhone or
+1.408.740.7256 (United States)
+1.888.240.2560 (US Toll Free)
+1.408.317.9253 (Alternate number)
(Global Numbers)
2) Enter Meeting ID: 543 903 672
3) Press #


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3540268
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20180228/71892cb6/attachment.html>

More information about the Hallc_running mailing list